Calculate the final temperature (in Kelvin) when 2.00 L at 294 K is compressed to 1.00 L. Your answer will be a whole number.
1
Expert's answer
2020-04-18T07:07:52-0400

V1/T1 = V2/T2

2.00 L / 294 K = 1.00 L / T2

T2 = 147 K